OpenStudy (anonymous):

Chan rows at a rate of 8 mph in still water. It takes him three hours to row upstream from his house to the park. He rows back home, and it takes him two hours. What is the speed of the current? Which of the following is a clue that would help solve this word problem? A. The rate of Chan’s rowing in still water is 8 mph. B. Chan rows from the house to the park. C. Chan rows back from the park to the house. D. The entire trip takes five hours.

OpenStudy (anonymous):

please help me!

OpenStudy (amistre64):

8 minus current, times time1 is distance traveled up 8 plus current, times time2 is distance traveled down when the distances equal .... we know the current 3(8-c) = 2(8+c) 24 -3c = 16 + 2c 24 - 16 = 5c c = (24-16)/5 so im not real sure what the extra conflagration is all about

OpenStudy (anonymous):

so would you think the answer would be C

OpenStudy (amistre64):

i dont know what the correct option would be. The information in the stated problem gives all the needed information ....

OpenStudy (amistre64):

A and D are evident, and B and C i spose are needed for clarification on mode of travel; but quite frankly none of them seem more right or wrong to me :/